An Olympic archer is able to hit the​ bull's-eye 84​% of the time. Assume each shot is independent of the others. She shoots 6 arrows.

​a) How many​ bull's-eyes do you expect her to​ get?

​b) What is the standard​ deviation?

​c) If she keeps shooting arrows until she hits the​ bull's-eye, how long do you expect it will​ take?
Answer #1

Let X be the number of 6 arrows to hit​ bull's-eye .

X ~ Binomial(n = 6, p = 0.84)

a)

Expected bull's-eyes to get hit = np = 6 * 0.84 = 5.04

b)

Standard​ deviation = \sqrt{np(1-p)} = \sqrt{6 * 0.84 * (1 - 0.84)}

= 0.898

c)

Let Y be the number shooting arrows until she hits the​ bull's-eye.

Y ~ Geometric(p = 0.84)

Expected number shooting arrows = E(Y) = 1/p

= 1/0.84

= 1.19